OpenStudy (anonymous):

\[\huge Q=Q_{0}e^{-\frac{t}{RC}}\] what would the charge be after10 time constants?

OpenStudy (jamesj):

when t = ...what, exactly?

OpenStudy (anonymous):

t=R*C

OpenStudy (jamesj):

If t = RC, then t/RC = 1 and hence \[ Q = Q_0 e^{-1} \] I think you might mean t = 10RC, in which case \[ Q = Q_0 e^{-10} \]

OpenStudy (jamesj):

e^10 = 22,026 approximately

OpenStudy (jamesj):

then either a) you are using a different time t from t = 10RC or b) you are calculating incorrectly

OpenStudy (jamesj):

...and I really with you'd stop deleting all of your replies! Anyway, I have to go. Food's getting cold.

OpenStudy (anonymous):

\[\huge Q=Q_0*5\times10^{-5}\]